Solve the equation 1 + 2log (x+1) = log (2x+1) = log (2x+1) +log (5x+8)


1
Expert's answer
2021-10-07T15:14:14-0400
"1 + 2\log (x+1) = \log (2x+1) +\log (5x+8)"

"\begin{matrix}
 x+1>0 \\
 2x+1>0 \\
 5x+8>0 
\end{matrix}=>x>-\dfrac{1}{2}"

Then


"\log\dfrac{(2x+1)(5x+8)}{(x+1)^2}=1"

"\dfrac{(2x+1)(5x+8)}{(x+1)^2}=10"

"10x^2+16x+5x+8=10x^2+20x+10"

"x=2"

"2>-\dfrac{1}{2}"

Answer: "x=2."